OpenStudy (anonymous):

Two factors of a polynomial, P(x), are (x + 6) and (x - 3). Which is a zero of P(x)? A. 6 B. -3 C. 3 D. 1

OpenStudy (anonymous):

3. p(x)=x-3 0=x-3 +3 +3 3=x.

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

No problem. :)